OpenStudy (anonymous):

complete the equation for the linear function whose graph contains the point of (5,-3) and has a slope of -1.25

OpenStudy (anonymous):

y=mx+b m=slope. y=-1.25x+b Plug the point into the equation -3=-1.25(5)+b solve for b -3=-6.25+b 3.25=b Plug them all into the equation, you have: y=-1.25x+3.25